# Feature Coding in the Era of Large Models: Dataset, Test Conditions, and Benchmark

### v3

@@ Abstract
Large models have achieved remarkable performance across various tasks, yet they incur significant computational costs and privacy concerns during both training and inference. Distributed deployment has emerged as a potential solution, but it necessitates the exchange of intermediate information between model segments, with feature representations serving as crucial information carriers. To optimize information exchange, feature coding is required to reduce transmission and storage overhead. Despite its importance, feature coding for large models remains an under-explored area. In this paper, we draw attention to large model feature coding and make three fundamental contributions. First, we introduce a comprehensive dataset encompassing diverse features generated by three representative types of large models. Second, we establish unified test conditions, enabling standardized evaluation pipelines and fair comparisons across future feature coding studies. Third, we introduce two baseline methods derived from widely used image coding techniques and benchmark their performance on the proposed dataset. These contributions aim to provide a foundation for future research and inspire broader engagement in this field. To support a long-term study, All source code and the dataset are made available at https://github.com/chansongoal/FCM-LM/tree/master.
@@ 1 Introduction
Existing feature coding research faces three key limitations: narrow model types, limited task diversity, and constrained data modalities. Most existing feature coding efforts focus on small-scale convolution-based networks, such as ResNet, ignoring large models. Although existing research emphasizes discriminative models, large models such as DINOv2 [31] and SAM [21] remain underrepresented. Generative models also have received little attention in this domain. Furthermore, existing research primarily targets visual features and neglects the investigation of textual features from large language models (LLMs). This gap indicates a need for broader exploration across both models and data modalities to advance feature coding.
In this work, we shift the feature coding research from small-scale to large-scale models. To address the existing limitations, we construct a comprehensive test dataset with features from five tasks across three large models, covering both visual and textual data for increased modality diversity. Furthermore, we establish unified test conditions to enable fair coding performance comparisons, including formulated bitrate computation and task-specific evaluation pipelines.
@@
  • We construct a comprehensive feature dataset encompassing three model types, five tasks, and two kinds of data modalities, along with unified test conditions to support long-term feature coding research.
  • We propose two baselines and evaluate their usability to the large model feature coding. We build a benchmark and the reported results suggest promising directions for further research.
@@ 2.1 Feature coding
Originally introduced within the Coding for Machines framework [45, 10], feature coding is one of two primary branches alongside visual coding. Unlike visual signal coding [36, 47, 15, 26, 37, 29], which encodes and reconstructs the original visual data, feature coding involves encoding features extracted from source data (e.g., images) into bitstreams. While visual coding has already been explored in the context of large models [38], feature coding remains underdeveloped in this domain.
Despite recent progress, feature coding research currently faces three key limitations. First, there is a restricted variety of model types in use. The existing feature extraction relies on small-scale convolution-based models such as ResNet series [24, 34, 20, 25, 3, 28, 17]. However, as Transformer architectures now dominate large model research [39, 11, 12, 31], it is crucial to investigate feature coding for Transformer-based models to capture the rich representations they provide. Second, the field is constrained by the limited scope of task types. Existing studies mainly address discriminative tasks like image classification and segmentation [7, 13, 48, 44, 5, 30], while generative tasks remain largely unexplored. Given the significance of generative models in modern AI, extending feature coding to include these tasks, such as image synthesis, is essential. The third limitation is the constrained range of source modality. Current feature coding research predominantly focuses on features extracted from visual data, overlooking the increasingly important features generated from textual data, particularly in light of advancements in LLMs. Expanding to multiple modalities is crucial as the diversity of AI applications grows.
Additionally, two core challenges hinder progress in feature coding for large models: the lack of a public test dataset and unified test conditions. Without a standardized dataset and evaluation conditions, fair comparisons across methods can not be conducted. A public dataset ensures consistent input data across studies, while unified test conditions establish a standard pipeline to evaluate performance.
These limitations motivate us to construct a diverse test dataset and formulate unified test conditions to advance feature coding research.
@@ 3.1.2 Cloud-edge distributed training
The cloud-edge distribution training pipeline is depicted in the middle of Fig. 1. The whole pipeline is divided into two parts: edge models and cloud models. During forward propagation, the edge model processes private data to generate features, which are then encoded into bitstreams and transmitted to the cloud. Upon receiving the bitstreams, the cloud decodes them back into features and feeds them into the cloud model. In the backward pass, gradients are computed and backpropagated along the reverse path. To optimize this process (reduce bandwidth costs and minimize latency), it is critical to apply feature coding before transmission.
@@ 3.2 Cloud-edge distributed inference
Once large models are trained, they are deployed in products to provide client services. For a product or service to succeed, two primary concerns must be addressed: supporting a high volume of simultaneous cloud server access and protecting user privacy. For the first concern, managing computational load is essential, particularly in generative applications that involve repetitive diffusion processes. A practical solution is to offload part of this computational demand to edge devices. For the second concern, the same strategy in the cloud-edge distributed training, as discussed in Sec. 3.1.2, can apply.
@@ 4.1.3 Source data collection
We noticed that learning-based methods may require a dedicated training set. We will provide the source code for generating features from any source data to serve as training data. Please find out our training data in the supplementary.
@@ 6 Baselines and benchmark
As a starting point, we introduce two baselines and establish a benchmark. We adapt these methods to account for the characteristics of large model features. Our goal is to examine their applicability to large model feature coding and provide insights for interested researchers, rather than to develop a completely new method.
@@ 7 Conclusion
Supplementary Material
Task Source Dataset Number of Samples Feature Shape Epoch Learning Rate Lambda
Cls ImageNet 5000 $256\times 256$ 800 1e-4 0.001, 0.0017, 0.003, 0.0035, 0.01
Seg VOC2012 5000 $256\times 256$ 800 1e-4 0.0005, 0.001, 0.003, 0.007, 0.015
Dpt NYUv2 20000 $256\times 256$ 200 1e-4 0.001, 0.005, 0.02, 0.05, 0.12
CSR Arc-Challenge, OpenBookQA 6000 $64\times 4096$ 200 1e-4 0.01405, 0.0142, 0.015, 0.15, 10
TTI Captions from COCO2017 50000 $512\times 512$ 60 1e-4 0.005, 0.01, 0.02, 0.05, 0.2
Table 7: Training details of the Hyperprior baseline.
8 Details of baseline methods
8.1 Packing details
8.1.1 Semantic segmentation
In semantic segmentation, an image is pre-processed into multiple $518\times 518$ patches by crop and resize operations. These patches are fed into DINOv2, which produces features of shape $1370\times 1536$. In the proposed dataset, each image is pre-processed into 2 patches, resulting in features with a shape of $2\times 1370\times 1536$. We reshape $F_{Seg}$ from $2\times 1370\times 1536$ into $2740\times 1536$ by vertically stacking them.
8.1.2 Depth estimation
In depth estimation, an image is first horizontally flipped. Both the original and flipped images are padded to dimensions that are multiples of 14 before being processed by DINOv2. For each image, DINOv2 generates $4\times 1611\times 1536$ features at the split point $SP_{DM}$. Thus, the shape of $F_{Dpt}$ is $2\times 4\times 1611\times 1536$. $F_{Dpt}$ is packed in two steps. First, each $4\times 1611\times 1536$ feature is horizontally packed into a $1611\times 6144$ feature by concatenating the features along the width. Second, the two $1611\times 6144$ features (original first and then flipped images) are vertically stacked, forming a $3222\times 6144$ feature.
8.1.3 Text-to-image synthesis
For text-to-image synthesis, the original feature $F_{TTI}$ has a shape of $16\times 128\times 128$, where 16 represents the number of channels. We divide the 16 channels into 4 subgroups. Within each subgroup, the $128\times 128$ features are horizontally concatenated to form a $128\times 512$ feature. Finally, the four $128\times 512$ features are vertically stacked to produce a final packed feature of size $512\times 512$.
8.2 Training details
8.2.1 Training dataset construction
We construct our training dataset from public datasets. The training features are extracted in the same way as the test dataset. The source code will be made publicly available, and we encourage researchers to extract features from their own source data.
As mentioned in the main text, raw features consume a significant amount of storage. During training, we encountered storage challenges. For example, 20000 original $F_{Dpt}$ features occupy 1474 GB of storage. As the demand for additional training data grows, storage requirements could increase substantially, often exceeding the resources available to researchers in academic institutions. To address this issue, we crop the features into smaller dimensions to facilitate feature coding training. The specific cropping configurations are detailed in Table 7.
It is worth emphasizing that while cropping is effective for feature coding training, this strategy is not applicable to large model training. For large models, cropping features to accelerate training is impractical, as downstream models typically require complete semantic information. For example, cropping a $256\times 256$ patch from $F_{Dpt}$ would make it challenging to perform accurate depth estimation. Therefore, designing efficient feature coding methods remains crucial for reducing the storage burden in large model deployments.
After feature extraction, the same truncation and quantization operations are applied to the training dataset. For the Hyperprior baseline, all features are normalized to the range [0, 1] before training.
8.2.2 Hyperparameter Setting
The hyperparameters, such as the number of training epochs, initial learning rate, and $\lambda$ values, are provided in Table 7. The number of training epochs is determined based on the size of the training dataset. In our experiments, all training processes converge to a stable loss. Five different $\lambda$ values are used to achieve various bitrates.
Metric QP Enc. Dec. QP Enc. Dec. QP Enc. Dec. QP Enc. Dec. QP Enc. Dec.
VTM Baseline 22 267.20 0.23 22 2301.30 1.11 22 14169.36 5.15 22 336.27 0.24 22 107.98 0.16
27 275.02 0.22 27 2039.34 0.94 27 10862.76 3.54 27 423.84 0.29 27 87.33 0.15
32 124.40 0.15 32 1399.71 0.52 32 7776.30 2.33 32 362.98 0.24 32 62.11 0.15
37 28.00 0.13 37 407.30 0.32 37 1590.27 1.14 37 162.05 0.17 37 32.38 0.24
42 8.55 0.14 42 70.66 0.27 42 348.68 1.17 42 66.28 0.18 42 15.70 0.15
Metric $\lambda$ Enc. Dec. $\lambda$ Enc. Dec. $\lambda$ Enc. Dec. $\lambda$ Enc. Dec. $\lambda$ Enc. Dec.
Hyperprior Baseline 0.01 0.04 0.06 0.015 0.3 0.38 0.12 1.57 2.05 10 0.07 0.12 0.2 0.03 0.03
0.0035 0.04 0.05 0.007 0.31 0.4 0.05 1.59 2.05 0.07 0.07 0.09 0.05 0.03 0.03
0.003 0.04 0.05 0.003 0.31 0.39 0.02 1.55 1.88 0.015 0.06 0.08 0.02 0.03 0.03
0.0017 0.05 0.06 0.001 0.29 0.35 0.005 1.54 1.84 0.0142 0.07 0.08 0.01 0.03 0.03
0.001 0.05 0.06 0.0005 0.29 0.36 0.001 1.57 1.91 0.01405 0.06 0.08 0.005 0.03 0.03
Table 8: Encoding and decoding complexity evaluations. “Enc.” and “Dec.” denote encoding time and decoding time, respectively. The running time is measured in seconds.
$\lambda$-BPFP
Semantic Segmentation
Common Sense Reasoning
Text-to-Image Synthesis
Figure 6: $\lambda$-BPFP curves and R-A curves of generalizability evaluation on Seg, CSR, and TTI tasks.
9 Additional experimental results
$SP_{DM1},QP=42$
$SP_{DM1},\lambda=0.001$
$SP_{DM2},QP=42$
$SP_{DM2},\lambda=0.001$
$SP_{DM3},QP=42$
$SP_{DM3},\lambda=0.001$
$SP_{DM4},QP=42$
$SP_{DM4},\lambda=0.001$
$SP_{GS},QP=32$
$SP_{GS},\lambda=0.01405$
$SP_{H},QP=42$
$SP_{H},\lambda=0.005$
Figure 7: Frequency and CDF visualization of the original truncated and reconstructed features. Green and red are for the original truncated features. Yellow and black are for the reconstructed features. Left: the VTM baseline. Right: the Hyperprior baseline.
9.1 Complexity analysis
We present the encoding and decoding complexity analyses in Table 8. The VTM baseline and Hyperprior baseline are evaluated on an Intel® Xeon® E5-2690 v4 CPU and a single NVIDIA GeForce RTX 4090 GPU, respectively. It is important to note that the VTM baseline is a handcrafted codec and cannot be executed on a GPU.
For the VTM baseline, encoding time increases as QP decreases and feature resolution increases. The block-based VTM codec exhaustively examines multiple partitions and modes during encoding, making it significantly more time-consuming than decoding. VTM is one of the reference implementations of the VVC standard. In practical deployments, a more efficient VVC codec, such as VVEnc, can be used to accelerate encoding. For the Hyperprior baseline, the decoding complexity is slightly higher than the encoding complexity. Across different bitrates, the encoding and decoding complexities remain comparable.
9.2 Generalizability analysis
Here, we visualize the $\lambda$-BPFP curves and the rate-accuracy curves for the cross evaluation to provide a more intuitive presentation of the results. The $\lambda$-BPFP curves reveal that the same $\lambda$ produces different rate-distortion trade-offs across various tasks, indicating that features extracted from different tasks do not share similar coding properties. From the rate-accuracy curves, we observe that coding models trained for one task struggle to achieve comparable performance when applied to another task. The generalizability of feature coding methods across multiple tasks has received limited attention in prior research. By presenting these results, we aim to highlight this important research direction and encourage further exploration.
9.3 Ablation on truncation and quantization
In this subsection, we examine the impact of truncation and quantization on feature coding. As shown in Table 10, truncation and quantization individually result in little task accuracy loss. When both are applied together, task accuracies experience slight decreases. However, applying the VTM baseline to quantized features leads to a significant reduction in accuracy.
Without truncation, feature values are often limited to a small subset within the full range of [0, 1023]. In this scenario, during feature coding, the codec’s internal quantization maps values with a small range to a single value, introducing significant distortion. By applying truncation, this range is expanded, which helps mitigate distortion caused by VTM’s internal quantization. To demonstrate this effect, we test several truncation ranges and encode the truncated features using fixed QPs (27, 32, 37). As presented in Table 11, smaller truncation ranges generally yield higher accuracy but come at the cost of increased bitrates. Therefore, selecting an appropriate truncation range is crucial for balancing the target bitrates and accuracy.
Proposed Features Existing Features
Source Dataset Split Point Max Min Mean IV GM Split Point Max Min Mean IV GM
ImageNet $SP_{DS}$ 79.01 -502.04 0.0782 17.04 3.97 $C5$-ReLU 7.81 0 0.4594 9839.49 143.22
VOC2012 $SP_{DS}$ 94.88 -485.23 0.0780 13.30 4.30 $C5$ 8.00 -4.24 0.2473 7323.72 131.90
NYUv2 $SP_{DM1}$ 3.25 -2.33 -0.0015 128.91 17.92 $p2$ 10.83 -$1$0.1726 8452.88 159.45
$SP_{DM2}$ 4.99 -26.08 -0.0005 36.96 10.00 $p3$ 8.35 -9.02 -0.1277 9418.53 168.38
$SP_{DM3}$ 23.51 -314.26 0.0042 9.61 2.76 $p4$ 9.06 -9.93 -0.0629 8499.73 158.89
$SP_{DM4}$ 96.04 -493.17 0.0731 12.54 4.35 $p5$ 9.61 -8.22 -0.0295 9402.06 169.13
Arc-Challenge $SP_{GS}$ 27.50 -26.38 0.0038 419.14 33.28 / / / / / /
Captions from COCO2017 $SP_{H}$ 3.67 -2.11 0.5664 21130.12 138.53 / / / / / /
Table 9: Statistics of the proposed features and existing features. (Gradient magnitude values are scaled by $10^{5}$).
Configuration Cls Seg Dpt CSR
Truncation Only 100 81.60 0.4992 100
Quantization Only 99 81.37 0.4982 100
Truncation + Quantization 100 79.93 0.4941 100
Quantization + VTM QP22 0 5.35 1.6085 6
Table 10: Ablations on truncation and quantization operations.
9.4 Distribution analysis on decoded features
We visualize the distribution of reconstructed features in Fig. 7, comparing the two baseline methods at similar bitrates. For $SP_{DM}$, the Hyperprior baseline fails to reconstruct features within their original distribution range. The reconstructed features appear truncated during the feature coding process. We attribute this truncation to the Hyperprior’s internal quantization (rounding operation), where the transformed latent variables are directly rounded to integers before entropy coding. In contrast, the VTM baseline quantizes original features into a broader distribution range of [0, 1023], effectively preventing the truncation phenomenon. A similar truncation phenomenon is observed for $SP_{H}$. For $SP_{GS}$, the features are encoded at a higher bitrate and the truncation phenomenon is not observed.
10 Additional dataset analysis
In addition to the visualizations provided in the main text, we present further statistical analyses of the proposed dataset.
10.1 Distribution analysis
We compare the statistical characteristics of the proposed features with those of existing features by randomly sampling 10 examples from each dataset and reporting their maximum, minimum, and mean values in Table 9. Overall, the proposed features exhibit a broader distribution range compared to existing features. In particular, for multi-scale features, the distribution range expands in deeper network layers, whereas existing features exhibit smaller variations. Moreover, although the proposed features tend to be more asymmetric, their mean values remain closer to zero.
Image Classification Semantic Segmentation
[-5, 5] [-30, 30] [-5, 5] [-30, 30]
BPFE Acc. BPFE Acc. BPFE mIoU BPFE mIoU
2.90 100 0.38 91 2.84 79.83 0.37 76.38
2.05 100 0.06 24 1.98 79.40 0.08 63.46
1.15 100 0.02 14 1.10 78.68 0.01 39.32
Depth Estimation Common Sense Reasoning
[-5, 5], [-5, 5] [-20, 20], [-30, 30] [-2, 2] [-10, 10]
BPFE RMSE BPFE RMSE BPFE Acc. BPFE Acc.
1.42 0.6490 0.58 0.8569 3.16 98 0.84 98
0.73 0.9604 0.17 1.1360 2.30 99 0.15 82
0.33 1.2453 0.04 1.3073 1.44 99 0.04 20
Table 11: Ablation results on truncation regions. $SP_{DM1}$ and $SP_{DM2}$ features are truncated into [-1, 1] and [-2, 2] in all cases, respectively.
10.2 Redundancy analysis
To evaluate spatial redundancy, we compute intensity variance (IV) and gradient magnitude (GM), where gradients are derived using the Sobel operator. Before these computations, both pixel and feature values are quantized to 10-bit integers. The intensity variance of the proposed features spans a wide range, from 9.61 to 22918.26, whereas existing features have a narrower range from 7323.72 to 9839.49. A similar trend is observed in gradient magnitude. These broader variations underscore the higher diversity of the proposed features. Furthermore, among the proposed features, textual features reveal distinct distribution characteristics compared to visual features. For instance, $SP_{GS}$ and $SP_{H}$ exhibit larger intensity variance and gradient magnitude. This observation highlights the necessity and importance of incorporating textual features in the dataset.
10.3 Visualization
We visualize the distributions, feature blocks, and DCT blocks of $F_{Cls}$ and $F_{Seg}$ extracted from $SP_{DS}$ in Fig. 8 and Fig. 9. These distributions exhibit high similarities to the features extracted from $SP_{DM4}$, as they are derived from the same layer. To avoid redundancy in the main paper, these visualizations have been moved here.
$SP_{DS},F_{Cls}$
$SP_{DS},F_{Seg}$
Figure 8: Visualization of the frequency distribution and CDF for the $F_{Cls}$ and $F_{Seg}$ features.
$SP_{DS},F_{Cls}$
$SP_{DS},F_{Seg}$
Figure 9: Visualization of the original feature blocks and their corresponding DCT blocks for the $F_{Cls}$ and $F_{Seg}$ features.

### v4

@@ Abstract
Large models have achieved remarkable performance across various tasks, yet they incur significant computational costs and privacy concerns during both training and inference. Distributed deployment has emerged as a potential solution, but it necessitates the exchange of intermediate information between model segments, with feature representations serving as crucial information carriers. To optimize information exchange, feature coding is required to reduce transmission and storage overhead. Despite its importance, feature coding for large models remains an under-explored area. In this paper, we draw attention to large model feature coding and make three fundamental contributions. First, we introduce a comprehensive dataset encompassing diverse features generated by three representative types of large models. Second, we establish unified test conditions, enabling standardized evaluation pipelines and fair comparisons across future feature coding studies. Third, we introduce two baseline methods derived from widely used image coding techniques and benchmark their performance on the proposed dataset. These contributions aim to provide a foundation for future research and inspire broader engagement in this field. To support a long-term study, all source code and the dataset are made available at https://github.com/chansongoal/LaMoFC.
footnotetext: This work was supported by the Ministry of Education of Singapore under Grant T2EP20123-0006. We acknowledge the support of GPU cluster built by MCC Lab of Information Science and Technology Institution, USTC.
@@ 1 Introduction
Existing feature coding research faces three key limitations: narrow model types, limited task diversity, and constrained data modalities. Most existing feature coding efforts focus on small-scale convolution-based networks, such as ResNet, ignoring large models. Furthermore, existing research primarily targets visual features and neglects the investigation of textual features from large language models (LLMs). This gap indicates a need for broader exploration across both models and data modalities to advance feature coding.
In this work, we shift the feature coding research from small-scale to large-scale models. To address the existing limitations, we construct a comprehensive test dataset with features from five tasks across three large models, covering both visual and textual data for increased modality diversity. Furthermore, we establish unified test conditions to enable fair comparisons, including formulated bitrate computation and task-specific evaluation pipelines.
@@
  • We construct a comprehensive feature dataset encompassing three model types, five tasks, and two kinds of source data modalities, along with unified test conditions to support long-term large model feature coding research.
  • We propose two baselines and evaluate their usability to the large model feature coding. We build a benchmark and suggest promising directions for further research.
@@ 2.1 Feature coding
Originally introduced within the Coding for Machines framework [48, 10], feature coding is one of two primary branches alongside visual coding. Unlike visual coding [39, 50, 15, 28, 40, 31, 35, 26, 25], which encodes and reconstructs the original visual data, feature coding involves encoding features extracted from source data into bitstreams. While visual coding has already been explored in the context of large models [41, 20], feature coding remains unexplored in this domain.
Despite recent progress, feature coding research still faces three key limitations. First, there is a restricted variety of model types in use. The existing feature coding focuses on small-scale CNN features such as [24, 37, 21, 27, 3, 30, 17]. However, as Transformer architectures now dominate large model research [42, 11, 12, 33], it is crucial to investigate feature coding for Transformer-based models. Second, the field is constrained by the limited scope of task types. Existing studies primarily address discriminative tasks such as image classification and segmentation [7, 13, 51, 47, 5, 32], while generative tasks remain largely unexplored. Given the significance of generative models, extending feature coding to include these tasks, such as image synthesis, is essential. The third limitation is the constrained range of source modality. Current feature coding research predominantly focuses on features extracted from visual data, overlooking the increasingly important features generated from textual data, particularly in light of advancements in LLMs. Expanding to multiple modalities is crucial as the diversity of AI applications grows.
Furthermore, two challenges hinder progress in large model feature coding: the lack of a public test dataset and unified test conditions. Without a standardized dataset and evaluation conditions, fair comparisons across methods can not be conducted. A public dataset ensures consistent input data across studies, while unified test conditions establish a standard pipeline to evaluate performance.
These limitations motivate us to construct a diverse test dataset and formulate unified test conditions to support large model feature coding research.
@@ 3.1.2 Cloud-edge distributed training
The cloud-edge distributed training pipeline is depicted in the middle of Fig. 1. The whole pipeline is divided into two parts: edge models and cloud models. During forward propagation, the edge model processes private data to generate features, which are then encoded into bitstreams and transmitted to the cloud. Upon receiving the bitstreams, the cloud decodes them back into features and feeds them into the cloud model. In the backward pass, gradients are computed and backpropagated along the reverse path. To reduce bandwidth cost and minimize latency, it is critical to apply feature coding before transmission.
@@ 3.2 Cloud-edge distributed inference
Once large models are trained, they are deployed in products to provide client services. For a product or service to succeed, two primary concerns must be addressed: supporting a high volume of simultaneous cloud server access and protecting user privacy. For the first concern, managing computational load is essential, particularly in generative applications that involve repetitive diffusion processes. A practical solution is to offload part of this computational demand to edge devices. For the second concern, the same strategy in the cloud-edge distributed training, as discussed in Sec. 3.1.2, can be applied.
@@ 4.1.3 Source data collection
We noticed that learning-based methods require a dedicated training set. Please follow the source code to generate training data and find our training data in the supplementary.
@@ 6 Baselines and benchmark
As a starting point, we introduce two baselines and establish a benchmark. We aim to examine their applicability to large model feature coding and provide insights for interested researchers.